OpenStudy (kainui):

Representing higher order derivatives in terms of first derivatives problems:

OpenStudy (kainui):

\[\Large f'(x) = \lim_{h \to 0} \frac{f(x+\frac{h}{2})-f(x-\frac{h}{2})}{h}\] This is my derivative definition, so if we apply it again we get the second derivative like this: \[\small f''(x) = \lim_{h \to 0} \frac{1}{h}(\frac{f(x+\frac{h}{2}+\frac{h}{2})-f(x+\frac{h}{2}-\frac{h}{2})}{h}-\frac{f(x-\frac{h}{2}+\frac{h}{2})-f(x-\frac{h}{2}-\frac{h}{2})}{h})\] simplify a little we get: \[f''(x) = \lim_{h \to 0} \frac{1}{h^2}[f(x+h)+f(x-h)-2f(x)]\] So with a little rearranging we can show an interesting fact that the concavity is proportional the the difference between the function at a point and the average of the two adjacent points, but that's not really what I'm looking for, I want to take the third derivative now and get a result, so we continue on and we can simplify to something that looks like this: \[\ f'''(x) = \lim_{h \to 0} \frac{1}{h^2}\frac{f(x+\frac{3h}{2})-f(x-\frac{3h}{2})}{h}-\frac{3}{h^2}\frac{f(x+\frac{h}{2})-f(x-\frac{h}{2})}{h}\]

OpenStudy (kainui):

Can we write this final thing in terms of just ordinary derivatives? It is tempting to just rewrite it by a change of variables, but I don't know if it's possible or how I would do it.

OpenStudy (anonymous):

If we apply the derivative a second time, I think we should end up with two limits.

OpenStudy (anonymous):

They may be equivalent, but I'm not completely sure just yet.

OpenStudy (kainui):

So instead of combining the h's use separate variables, h, then k, etc? I am not sure either but it is an idea to play with.

OpenStudy (anonymous):

Start like this, then use definition of derivatives:\[ \lim_{k\to 0}\frac{f'(x+k)-f'(x)}{k} \]

OpenStudy (kainui):

By changing variables and separating the limit into two parts we can have: \[\ f'''(x) = 216\lim_{h \to 0} \frac{1}{h^2}\frac{f(x+\frac{h}{2})-f(x-\frac{h}{2})}{h}-3\lim_{h \to 0}\frac{1}{h^2}\frac{f(x+\frac{h}{2})-f(x-\frac{h}{2})}{h}\] however I don't even know if I'm technically allowed to separate this limit or not.

OpenStudy (kainui):

I think that will just give essentially the same thing @wio I just kind of like looking from this perspective since I can sort of see symmetry a little easier.

OpenStudy (anonymous):

Well for one, does the order of taking limits matter? You're assuming it doesn't when you merge them.

OpenStudy (kainui):

Order of taking limits? Is it even possible to take the second derivative before the first derivative?

OpenStudy (anonymous):

I mean letting \(k\) goto \(0\) before \(h\) dies.

OpenStudy (kainui):

Well what is the difference between k and h?

OpenStudy (kainui):

If I switched them when you weren't looking you wouldn't tell. It's like if I have x and want to add 2. I can either add 1 first and then add 1 or I can add 1 then add 1 the other way around, and no one can tell the difference I think.

OpenStudy (anonymous):

Start like this, then use definition of derivatives:\[ \lim_{k\to 0}\frac{f'(x+k)-f'(x)}{k} =\lim_{k\to 0}\lim_{h\to 0}\frac{f(x+h+k)-f(x+k)-f(x+h)+f(x)}{hk} \]

OpenStudy (kainui):

Sure, now replace h and k with each other, they are completely symmetric: \[ \lim_{k\to 0}\lim_{h\to 0}\frac{f(x+h+k)-f(x+k)-f(x+h)+f(x)}{hk} = \ \ \]\[ \lim_{h\to 0}\lim_{k\to 0}\frac{f(x+k+h)-f(x+h)-f(x+k)+f(x)}{kh} \]

OpenStudy (anonymous):

So if they're symmetric then if follows that we can let \(h=k\) I suppose?

OpenStudy (kainui):

I think so, since I can imagine bringing h to 0 is the same as making k to 0 so in the hk-plane we should be able to just go straight to 0 in a single line 

OpenStudy (kainui):

I have seen it done before by a differential equations book I owned, but then again just because it works doesn't necessarily mean anything.

OpenStudy (anonymous):

Were aren't taking them at the same time, we're taking them in order, which does make things easier.

OpenStudy (anonymous):

We've already assumed that \(f'(x)\) exists.

OpenStudy (kainui):

Yeah this is all done based on the idea that all the higher derivatives we need exist, that's not important to worry about.

OpenStudy (anonymous):

I think since we assume the derivatives exist, then we can use: \[ \lim_{x\to a}g(x) = b\land \lim_{x\to b}f(x) = L \implies \lim_{x\to a}f(g(x)) = L \]

OpenStudy (anonymous):

In the case the function behaves relatively nicely, the two-limits of a second order derivative can be reduced to a single limit--but keep in mind this is the exception, not the rule. $$f''(x)=\frac{df'}{dx}=\lim_{\delta x\to0}\frac{f'(x+\delta x)-f'(x)}{\delta x}$$ ... is the general definition of the second derivative of a differentiable function \(f:\mathbb{R}\to\mathbb{R}\). If you're interested in similar limits for higher order derivatives then you should check out finite differences. http://en.wikipedia.org/wiki/Finite_difference

OpenStudy (anonymous):

http://en.wikipedia.org/wiki/Finite_difference#Higher-order_differences In the limit as the step size or spacing goes to 0, the general formulae for the higher-order finite differences agree with the corresponding derivatives (for sufficiently well-behaved functions).

OpenStudy (anonymous):

Okay, I'm curious as to what the exceptions are.

OpenStudy (anonymous):

well behaved means \(C^1\)?

OpenStudy (anonymous):

I think the exceptions would fall along the lines of \(f(x)\) is not differentiable and yet the higher order difference limit just so happens to exist.

OpenStudy (anonymous):

I honestly didn't think too deeply about any particular conditions for 'well-behaved' but I figured I'd play it safe :-p I'll give it some thought.

OpenStudy (anonymous):

but indeed, you can find more of these single-limit formulae using the n-th order finite differences and simply taking their limit as the step size goes to 0.

OpenStudy (anonymous):

You're right. Proving anything complicated concerning limits ends but being a major battle.

OpenStudy (anonymous):

Differentiable is a very high criterion, typically speaking.